You have written a call option on Walmart common stock. The option has an exercise price of $87, and Walmart’s stock currently trades at $85. The option premium is $1.25 per contract.

a. How much of the option premium is due to intrinsic value versus time value?
b. What is your net profit if Walmart’s stock price decreases to $83 and stays there until the option expires?
c. What is your net profit on the option if Walmart’s stock price increases to $93 at expiration of the option and the option holder exercises the option?
Answer #1

a. Intrinsic value = 0 because Current stock price < Exercise price

Time value = entire premium = $1.25

b. Profit of a short call = -max(St - X, 0) + premium received

Profit = -max(83 - 87, 0) + 1.25

Profit = $1.25

c. Profit = -max(93 - 87, 0) + 1.25

Profit = -max(6,0) + 1.25

Profit = -6 + 1.25

Profit = -$4.75

Can you please upvote? Thank You :-)